# Training Robust Graph Neural Networks by Modeling Noise Dependencies

### v1

@@ 1 Introduction
In recent years, graph neural networks (GNNs) have demonstrated remarkable achievements in graph representation learning and have been extensively applied in numerous downstream tasks (Yao et al., 2019; Wang et al., 2019; Kim et al., 2023, 2024). However, in the majority of real-world scenarios, node features frequently exhibit noise due to various factors, leading to the creation of inaccurate graph representations (Liu et al., 2021; Jin et al., 2022). For instance, in user-item graphs, users may create fake profiles or posts, and fraudsters and malicious users may write fake reviews or content on items, resulting in noisy node features. Recent studies have revealed the vulnerability of GNNs to such scenarios, highlighting the necessity to design robust GNN models against noisy node features.
Figure 1: Examples of DANG in social networks: IFN represents independent node feature noise. Under the IFN assumption (b), Bob’s noisy features are presumed to have no effect on the graph structure or node labels. However, in real-world DANG scenarios (c), Bob’s noisy features can propagate, leading to both structural noise in the graph and label noise.
@@
To this end, various methods have been proposed to make a huge success in terms of model robustness (Liu et al., 2021; Jin et al., 2022). These methods are founded on the independent node feature noise (IFN) assumption, which posits that noise in node features does not impact the graph structure or node labels. Under the IFN assumption (Fig. 1(b)), for example, Bob’s fake profile does not influence other nodes, which is also explained by the data generating process (DGP) of IFN (See Fig. 2(a)) in which no causal relationships exist among the noisy node features $X$, graph structure $A$, and node labels $Y$.
@@
However, we should rethink: In real-world applications, can noise in node features truly be isolated from influencing the graph structure or node labels? Let us explore this through examples from social networks (Fig. 1). Consider Bob, who introduces noisy node features by creating fake profiles or posts. Other users, such as Alice and Tom, may then connect with Bob based on his fake profile, resulting in noisy connections that contribute to graph structure noise. Over time, these noises could alter the community associations of Alice and Tom, leading to noisy node labels. Such causal relationships among $X$, $A$, and $Y$ (i.e., $A\leftarrow X$, $Y\leftarrow X$, and $Y\leftarrow A$) are depicted in Fig. 2(b). Similarly, in an e-commerce network, malicious users may post fake reviews or content about items, creating noisy node features. These fake interactions could lead other users to engage with irrelevant items, introducing noisy co-interaction edges between items (i.e., graph structure noise). As a result, automated item category labeling systems, which rely on both node features and graph structure, might inaccurately label item categories, resulting in label noise.
@@
These scenarios underscore an important insight: In real-world applications, noise in node features may create a chain of noise dependencies that propagate to the graph structure and node labels. This highlights the pressing need for robust GNNs capable of addressing such noise dependencies — an aspect that has been largely overlooked in current research. Since such noise dependencies are prevalent across a wide range of real-world applications [^1] Additional real-world examples demonstrating the practical existence of such noise are provided in Sec 3., failing to address them can result in significant robustness gaps and impede the development of more practical and robust GNN models. However, we observe that existing robust GNN models indeed fail to generalize effectively in such noise scenario since they overlook the underlying relationships among $X$, $A$, and $Y$ within the data generation process.
@@
To enhance the practicality of existing noise assumptions and robust GNNs, we newly introduce a novel dependency-aware noise on graphs (DANG) and propose a dependency-aware robust graph neural network framework (Da-Gnn) that directly models the DGP of DANG. We first illustrate the DGP of DANG as shown in Fig. 2(b) (c.f. Sec 3). More precisely, we introduce three observable variables (i.e., $X$, $A$, and $Y$) and three latent variables (i.e., noise incurring variable $\epsilon$, latent clean graph structure $Z_{A}$, and latent clean node labels $Z_{Y}$), while defining causal relationships among these variables to represent the data generation process of DANG. We then devise a deep generative model, Da-Gnn, that directly captures the causal relationships among the variables in the DGP of DANG by 1) deriving a tractable learning objective based on variational inference (c.f. Sec 4.1) and 2) addressing non-trivial technical challenges in implementing the learning objective (c.f. Sec 4.2). Moreover, to rigorously evaluate our proposed method, we propose both synthetic and real-world DANG benchmark datasets. In our experiments, we demonstrate that Da-Gnn effectively generalizes not only to DANG but also to other noise assumptions commonly considered in this field of research. This highlights Da-Gnn’s broader applicability compared to existing robust GNN models. In summary, the main contributions of our paper are as follows:
  • To the best of our knowledge, this is the first study to examine the gap between real-world scenarios and the overly simplistic noise assumptions underlying previous robust GNN research, which constrain their practicality.
@@
  • We propose novel graph benchmark datasets that simulate DANG within real-world applications. To the best of our knowledge, this is the first work proposing new datasets for evaluating the robust GNNs under realistic noise scenarios that are plausible in real-world applications, which fosters practical research in robust GNNs.
@@ 2.1 Noise-Robust GNN
Feature noise-robust GNN. AirGNN (Liu et al., 2021) identifies and addresses nodes with noisy features based on the hypothesizes that they tend to have dissimilar features within their local neighborhoods. Consequently, this approach tackles the noisy node features while assuming that the structure of the input graph is noise-free.
Structure noise-robust GNN. RSGNN (Dai et al., 2022) aims to train a graph structure learner by encouraging the nodes with similar features to be connected. STABLE (Li et al., 2022a) removes edges with low feature similarity, learns node representations from the modified structure, and constructs a kNN graph as the refined structure. In summary, these methods tackle the noisy graph structure while assuming that node features are noise-free.
Label noise-robust GNN. NRGNN (Dai et al., 2021) connects edges between nodes with similar features, mitigating the information propagation from falsely labeled nodes. RTGNN (Qian et al., 2023) uses small-loss approach (Han et al., 2018), but nodes with noisy features or structures exhibit large losses, leading to inaccuracies of the approach. Therefore, these methods tackle the noisy node labels while assuming that both node features and graph structure are noise-free.
Multifaceted noise-robust GNN. SG-GSR (In et al., 2024) tackles multifaceted structure and feature noise by identifying a clean subgraph within a noisy graph structure and augmenting it using label information. This augmented subgraph serves as supervision for robust graph structure refinement. However, since noisy label information can compromise the augmentation process, SG-GSR relies on the assumption that node labels are free of noise.
@@ 2.2 Generative Approach
Yao et al. (2021) devises a generative approach to model the DGP of instance-dependent label noise (Berthon et al., 2021). However, extending this method to the graph domain introduces significant challenges. It requires handling additional latent variables and complex causal relationships, such as $Z_{A}$, $\epsilon_{A}$, $A\leftarrow\epsilon_{A}$, $A\leftarrow X$, $Y\leftarrow A$, and $A\leftarrow Z_{A}$, each posing non-trivial obstacles beyond the straightforward extension [^2] Detailed explanation is outlined in Appendix D.1. WSGNN (Lao et al., 2022) and GraphGLOW (Zhao et al., 2023) utilize a probabilistic generative approach and variational inference to infer the latent graph structure and node labels. However, they assume noise-free graphs, reducing effectiveness in real-world noisy scenarios.
@@ 3 Dependency-Aware Noise on Graphs
Formulation. In this section, we define a new graph noise assumption, DANG, and its DGP. In Fig. 2(b), $X$ denotes the node features (potentially noisy), $Y$ the observed node labels (possibly noisy), $A$ the observed edges (which may contain noise), and $\epsilon$ the environment variable causing the noise. $Z_{Y}$ represents the latent clean node labels, while $Z_{A}$ does the latent clean graph structure encompassing all potential node connections. We give the explanations for each causal relationship within the DGP of DANG along with the examples in user graphs in social networks:
@@
$A\leftarrow\epsilon$: To provide a broader scope, we also posit that $\epsilon$ is a potential cause of $A$. This extension is well-founded, as real-world applications often exhibit graph structure noise originating from various sources in addition to the feature-dependent noise (Liu et al., 2022; Fatemi et al., 2021).
$Y\leftarrow(Z_{Y},X,A)$: $Z_{Y}$, $X$, and $A$ are causes of $Y$. To give an example, the true communities (or interests) of users (i.e., $Z_{Y}$) are leveraged to promote items to targeted users within a community (Ma et al., 2021). To detect the communities, both node features and graph structures are utilized. However, if a user has noisy node features (i.e., $X$) or noisy edges (i.e., $A$), the user may be assigned to a wrong community (or interest), which leads to noisy labels (i.e., $Y$).
For simplicity, we assume $\epsilon$ is not a cause of $Y$. This assumption matches real-world scenarios where mislabeling is more likely due to confusing or noisy features rather than arbitrary sources (Berthon et al., 2021). In other words, label noise in graphs is predominantly caused by confusing or noisy features and graph structure (i.e., $Y\leftarrow(X,A)$), rather than an arbitrary external factor (i.e., $Y\nleftarrow\epsilon$).
Discussion. 1) Under DANG a graph does not contain any noise-free data sources. This point presents a non-trivial challenge for the existing robust GNN methods to tackle DANG, as they assume the completeness of at least one data source.
2) DANG is prevalent across various domains, including social networks, e-commerce, web graphs, citation networks, and biology. Due to space constraints, detailed scenarios are provided in Appendix C.1. While not all noise scenarios perfectly align with DANG, such cases are rare compared to the widespread occurrence of DANG in these domains, dominating graph applications.
@@ 4.1 Problem Formulation
Learning Objective. We adopt the variational inference framework (Blei et al., 2017; Yao et al., 2021) to optimize the Evidence Lower-BOund (ELBO) of the marginal likelihood for observed data, i.e., $P(X,A,Y)$, rather than optimizing the marginal likelihood directly. Specifically, we derive the negative ELBO, i.e., $\mathcal{L}_{\text{ELBO}}$, as follows:
@@
where $kl(\cdot||\cdot)$ denotes KL divergence. The derivation details are provided in Appendix A. The encoders $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$ infer three latent variables, while $\theta_{1}$, $\theta_{2}$, and $\theta_{3}$ are decoder networks used for generating three observable variables. Our objective is to find the optimal values of ϕ=italic-$\phi=$$\phi_{1}$, $\phi_{2}$, $\phi_{3}$} and $\theta=$$\theta_{1}$, $\theta_{2}$, $\theta_{3}$} that minimize the value of $\mathcal{L}_{\text{ELBO}}$. By doing so, the encoders and decoders are trained to directly capture the causal relationships among the variables that introduce noise. Consequently, it promotes the accurate inference of the latent clean node label $Z_{Y}$ and latent clean graph structure $Z_{A}$ to effectively perform the node classification and link prediction tasks even in the presence of DANG.
@@ 4.2 Model Instantiations
In this section, we present the details of the practical implementation and optimization of Da-Gnn based on the learning objective, $\mathcal{L}_{\text{ELBO}}$. The overall architecture and detailed algorithm of Da-Gnn are provided in Fig 3 and Algorithm 1 in Appendix, respectively. The key challenge of the instantiation is how to accurately infer the latent variables $Z_{A}$, $Z_{Y}$, and $\epsilon$ in the presence of noisy $X$, $A$, and $Y$. To alleviate the challenge, we design the robust inference encoders (Fig 3(a)) and generative decoders (Fig 3(b)) with the corresponding regularizers (Fig 3(a)) and reconstruction losses (Fig 3(b)). Consequently, the encoders would be able to accurately infer the latent variables by capturing the causal relationships among the variables that introduce noise.
@@ 4.2.1 Modeling Inference Encoder
Modeling $q_{\phi_{1}}(Z_{A}|X,A)$. The objective of modeling $q_{\phi_{1}}(Z_{A}|X,A)$ is to accurately infer the latent clean graph structure $Z_{A}$ that enhances the message passing of a GNN model. To this end, we design the encoder $\phi_{1}$ as a graph structure learner. More specifically, to obtain the latent graph $\mathbf{\hat{A}}=\{\hat{a}_{ij}\}_{N\times N}$, we sample from $\hat{a}_{ij}\sim\text{Bern}(\hat{p}_{ij})$ where $\hat{p}_{ij}=\rho(s(\mathbf{Z}_{i},\mathbf{Z}_{j}))$ and $\mathbf{Z}=\text{GCN}_{\phi_{1}}(\mathbf{X},\mathbf{A})$. To further promote an accurate inference of $Z_{A}$, we regularize $Z_{A}$ to mainly include assortative edges (Zhao et al., 2023) by minimizing $kl(\hat{\mathbf{A}}||\mathbf{A}^{\text{p}})$, where $\mathbf{A}^{\text{p}}=\{{a}^{\text{p}}_{ij}\}_{N\times N}$ and ${a}^{\text{p}}_{ij}$ is sampled from a Bernoulli distribution with a probability as given by the $\gamma$-hop subgraph similarity (Choi et al., 2022; Dai et al., 2022). It is important to note that this regularization is equivalent to minimizing $kl(q_{\phi_{1}}(Z_{A}|X,A)||p(Z_{A}))$ in Eqn. 1. However, computing $\hat{p}_{ij}$ in every epoch is impractical for large graphs, i.e., $O(N^{2})$. To this end, we pre-define a proxy graph based on the subgraph similarity, and compute $\hat{p}_{ij}$ as edge weights on the proxy graph. Please refer to the Appendix B for detailed information on implementation details.
Modeling $q_{\phi_{3}}(Z_{Y}|X,A)$. The objective of modeling $q_{\phi_{3}}(Z_{Y}|X,A)$ is to accurately infer the latent clean node label $Z_{Y}$. To this end, we instantiate the encoder $\phi_{3}$ as a GCN classifier. Specifically, we infer $Z_{Y}$ through $\hat{\mathbf{Y}}=\text{GCN}_{\phi_{3}}(\mathbf{X},\mathbf{\hat{A}})\in\mathbb{R}^{N\times C}$. We introduce the node classification loss ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{L}_{\text{cls-enc}}}=\sum_{i\in\mathcal{V}^{L}}\text{CE}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{i})$, where CE is the cross entropy loss. To further enhance the quality of inference of $Z_{Y}$, we regularize $Z_{Y}$ to satisfy class homophily (McPherson et al., 2001) by minimizing the KL divergence between the probability predictions $\mathbf{\hat{Y}}$ of each node and its first order neighbors in $\mathbf{\hat{A}}$. The implemented loss function is given by: ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{L}_{\text{hom}}}=\sum_{i\in\mathcal{V}}\frac{\sum_{j\in\mathcal{N}_{i}}\hat{p}_{ij}\cdot kl(\hat{\mathbf{Y}}_{j}||\hat{\mathbf{Y}}_{i})}{\sum_{j\in\mathcal{N}_{i}}\hat{p}_{ij}}$, where $\mathcal{N}_{i}$ denotes the set of first-order neighbors of node $v_{i}$ within $\hat{\mathbf{A}}$. It is worth noting that this regularization is equivalent to minimizing $kl(q_{\phi_{3}}(Z_{Y}|X,A)||p(Z_{Y}))$ in Eqn. 1.
@@
The objective of modeling $q_{\phi_{22}}(\epsilon_{A}|X,A)$ is to accurately infer the structure noise incurring variable $\epsilon_{A}$ that determines whether each edge is clean or noisy. To this end, we regard $\epsilon_{A}$ as a set of scores indicating the likelihood of each observed edge being clean or noisy. To estimate the likelihood, we utilize small loss approach (Han et al., 2018). Precisely, we compute the set of link prediction losses as $\{(1-\hat{p}^{el}_{ij})^{2}|(i,j)\in\mathcal{E}\}$, where $\hat{p}^{el}_{ij}$ represents the $\hat{p}_{ij}$ value at the final epoch during early-learning phase. Therefore, an edge with high $\hat{p}^{el}_{ij}$ value can be considered as a clean edge, and we instantiate $\epsilon_{A}$ as $\{\hat{p}_{ij}^{el}|(i,j)\in\mathcal{E}\}$. To alleviate the uncertainty of a single training point’s loss value, we adopt an exponential moving average (EMA) technique: $\hat{p}_{ij}^{el}\leftarrow\xi\hat{p}_{ij}^{el}+(1-\xi)\hat{p}_{ij}^{c}$, where $\hat{p}_{ij}^{c}$ indicates the value of $\hat{p}_{ij}$ at the current training point, and $\xi$ indicates the decaying coefficient fixed to 0.9. This approach is equivalent to minimizing $kl(q_{\phi_{22}}(\epsilon_{A}|X,A)|p(\epsilon_{A}))$, where $p(\epsilon_{A})$ is assumed to follow the same distribution as $q_{\phi_{22}}(\epsilon_{A}|X,A)$ but with lower variance.
For the encoder $\phi_{21}$, we use an MLP that takes $X$ and $Z_{Y}$ as inputs and infers $\epsilon_{X}$. Additionally, we regularize $p(\epsilon_{X})$ to follow the standard multivariate normal distribution, which means that a closed form solution of $kl(q_{\phi_{21}}(\epsilon_{X}|X,Z_{Y})||p(\epsilon_{X}))$ can be obtained as ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{L}_{\text{p}}}=-\frac{1}{2}\sum_{j=1}^{d_{2}}(1+\log\sigma_{j}^{2}-\mu_{j}^{2}-\sigma_{j}^{2})$ (Kingma & Welling, 2013), where $d_{2}$ is the dimension of a $\epsilon_{X}$. Note that these two regularization techniques are equivalent to minimizing $\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[kl(q_{\phi_{2}}(\epsilon|X,A,Z_{Y})||p(\epsilon))\right]$ in Eqn. 1.

### v2

@@ 1 Introduction
In recent years, graph neural networks (GNNs) have demonstrated remarkable achievements in graph representation learning and have been extensively applied in numerous downstream tasks gcntext ; ngcf ; kim2023class ; kim2024revisiting . However, in the majority of real-world scenarios, node features frequently exhibit noise due to various factors, leading to the creation of inaccurate graph representations airgnn ; jin2022empowering . For instance, in user-item graphs, users may create fake profiles or posts, and fraudsters and malicious users may write fake reviews or content on items, resulting in noisy node features. Recent studies have revealed the vulnerability of GNNs to such scenarios, highlighting the necessity to design robust GNN models against noisy node features.
@@
To this end, various methods have been proposed to make a huge success in terms of model robustness airgnn ; jin2022empowering . These methods are founded on the independent node feature noise (IFN) assumption, which posits that noise in node features does not impact the graph structure or node labels. Under the IFN assumption (Fig. 1(b)), for example, Bob’s fake profile does not influence other nodes, which is also explained by the data generating process (DGP) of IFN (See Fig. 2(a)) in which no causal relationships exist among the noisy node features $X$, graph structure $A$, and node labels $Y$.
@@
However, we should rethink: In real-world applications, can noise in node features truly be isolated from influencing the graph structure or node labels? Let us explore this through examples from social networks (Fig. 1). Consider Bob, who introduces noisy node features by creating fake profiles or posts. Other users, such as Alice and Tom, may then connect with Bob based on his fake profile, resulting in noisy connections that contribute to graph structure noise. Over time, these noises could alter the community associations of Alice and Tom, leading to noisy node labels. Such causal relationships among node features $X$, graph structure $A$, and node label $Y$ (i.e., $A\leftarrow X$, $Y\leftarrow X$, and $Y\leftarrow A$) are depicted in Fig. 2(b).
Figure 1: Examples of DANG in social networks: IFN represents independent node feature noise. Under the IFN (b), Bob’s noisy features have no effect on the graph structure or node labels. However, in DANG (c), Bob’s noisy features can propagate, leading to both structural noise in the graph and label noise.
@@
This scenario underscore an important insight: In real-world applications, noise in node features may create a chain of noise dependencies that propagate to the graph structure and node labels. This highlights the pressing need for robust GNNs capable of addressing such noise dependencies, an aspect that has been largely overlooked in current research.
@@
Since such noise dependencies are prevalent across a wide range of real-world applications [^1] Additional real-world examples demonstrating the practical existence of such noise are provided in Sec 3. in addition to social networks, failing to address them can result in significant robustness gaps and impede the development of more practical and robust GNN models. However, we observe that existing robust GNN models indeed fail to generalize effectively in such noise scenario since they overlook the underlying relationships among $X$, AA, and $Y$ within the data generation process.
To enhance the practicality of existing noise assumptions and robust GNNs, we newly introduce a dependency-aware noise on graphs (DANG) and propose a dependency-aware robust graph neural network framework (DA-GNN) that directly models the DGP of DANG. We first illustrate the DGP of DANG as shown in Fig. 2(b) (c.f. Sec 3). More precisely, we introduce three observable variables (i.e., $X$, AA, and $Y$) and three latent variables (i.e., noise incurring variable $\epsilon$, latent clean graph structure $Z_{A}$, and latent clean node labels $Z_{Y}$), while defining causal relationships among these variables to represent the data generation process of DANG. We then devise a deep generative model, DA-GNN, that directly captures the causal relationships among the variables in the DGP of DANG by 1) deriving a tractable learning objective based on variational inference (c.f. Sec 4.1) and 2) addressing non-trivial technical challenges in implementing the learning objective (c.f. Sec 4.2). Moreover, to rigorously evaluate our proposed method, we propose both synthetic and real-world DANG benchmark datasets. In our experiments, we demonstrate that DA-GNN effectively generalizes not only to DANG but also to other noise assumptions commonly considered in this field of research. This highlights DA-GNN’s broader applicability compared to existing robust GNN models. In summary, the main contributions of our paper are as follows:
  • We examine the gap between real-world scenarios and the overly simplistic noise assumptions underlying previous robust GNN research, which constrain their practicality.
@@
  • We propose novel graph benchmark datasets that simulate DANG in real-world applications to evaluate robust GNNs under realistic and plausible noise conditions, thereby promoting practical research in robust graph learning.
@@ 2.1 Noise-Robust GNN
Feature noise-robust GNN. AirGNN airgnn identifies and addresses nodes with noisy features based on the hypothesis that they tend to have dissimilar features within their local neighborhoods. Consequently, this approach tackles the noisy node features while assuming that the structure of the input graph is noise-free.
Structure noise-robust GNN. RSGNN rsgnn aims to train a graph structure learner by encouraging the nodes with similar features to be connected. STABLE stable removes edges with low feature similarity, learns node representations from the modified structure, and constructs a kNN graph as the refined structure. In summary, these methods tackle the noisy graph structure while assuming that node features are noise-free.
Label noise-robust GNN. Although there have been many label noise-robust GNNs nrgnn ; rtgnn ; wu2024mitigating ; pignn ; erase ; graphcleaner ; ding2024divide , all these methods are built on the assumption that either node features or graph structures are noise-free. For example, RTGNN rtgnn uses small-loss approach han2018co , but nodes with noisy features or structures exhibit large losses, leading to inaccuracies of the approach. TSS wu2024mitigating mitigates label noise relying on the structural information, which can be noisy.
Multifaceted noise-robust GNN. SG-GSR in2024self tackles multifaceted structure and feature noise by identifying a clean subgraph within a noisy graph structure and augmenting it using label information. This augmented subgraph serves as supervision for robust graph structure refinement. However, since noisy label information can compromise the augmentation process, SG-GSR relies on the assumption that node labels are free of noise.
@@ 2.2 Generative Approach
yao2021instance devises a generative approach to model the DGP of instance-dependent label noise berthon2021confidence . However, extending this method to the graph domain introduces significant challenges. It requires handling additional latent variables and complex causal relationships, such as $Z_{A}$, $\epsilon_{A}$, $A\leftarrow\epsilon_{A}$, $A\leftarrow X$, $Y\leftarrow A$, and $A\leftarrow Z_{A}$, each posing non-trivial obstacles beyond the straightforward extension [^2] Detailed explanation is outlined in Appendix D.. WSGNN lao2022variational and GraphGLOW zhao2023graphglow utilize a probabilistic generative approach and variational inference to infer the latent graph structure and node labels. However, they assume noise-free graphs, reducing effectiveness in real-world noisy scenarios.
@@ 3 Dependency-Aware Noise on Graphs
3.1 Formulation
In this section, we define a new graph noise assumption, DANG, and its DGP. In Fig. 2(b), $X$ denotes the node features (potentially noisy), $Y$ denotes the observed node labels (possibly noisy), $A$ denotes the observed edges (which may contain noise), and $\epsilon$ denotes the environment variable causing the noise. $Z_{Y}$ represents the latent clean node labels, while $Z_{A}$ does the latent clean graph structure encompassing all potential node connections. We give the explanations for each causal relationship within the DGP of DANG along with the examples in user graphs in social networks:
@@ 3.1 Formulation
  • $A\leftarrow\epsilon$: To provide a broader scope, we also posit that $\epsilon$ is a potential cause of $A$. This extension is well-founded cogsl ; slaps , as real-world applications often exhibit graph structure noise originating from various sources in addition to the feature-dependent noise.
  • $Y\leftarrow(Z_{Y},X,A)$: $Z_{Y}$, $X$, and $A$ are causes of $Y$. To give an example, the true communities (or interests) of users (i.e., $Z_{Y}$) are leveraged to promote items to targeted users within a community communitysurvey . To detect the communities, both node features and graph structures are utilized. However, if a user has noisy node features (i.e., $X$) or noisy edges (i.e., $A$), the user may be assigned to a wrong community (or interest), which leads to noisy labels (i.e., $Y$).
For simplicity, we assume $\epsilon$ is not a cause of $Y$. This assumption matches real-world scenarios where mislabeling is more likely due to confusing or noisy features rather than arbitrary sources berthon2021confidence . In other words, label noise in graphs is predominantly caused by confusing or noisy features and graph structure (i.e., $Y\leftarrow(X,A)$), rather than an arbitrary external factor (i.e., $Y\nleftarrow\epsilon$).
3.2 Discussion
1) Under DANG a graph does not contain any noise-free data sources. This point presents a non-trivial challenge for the existing robust GNN methods to tackle DANG, as they assume the completeness of at least one data source.
2) DANG is prevalent across diverse domains, including social, e-commerce, web, and biological graphs. Due to space constraints, detailed statistical evidences and intuitive examples on the existence of DANG in real-world applications are provided in Appendix C.1 and C.2. We acknowledge, however, that not all noise scenarios perfectly align with DANG. For instance, in non-relational domains such as molecular structures or protein–protein interaction networks, the graph structure is fixed and unaffected by node feature noise. Nevertheless, we claim that such cases are rare compared to the broad applicability of DANG across widely studied graph domains, including social, e-commerce, web, and biological (cell-cell) networks.
@@ 4.1 Problem Formulation
Learning Objective. We adopt the variational inference framework vi_review ; yao2021instance to optimize the Evidence Lower-BOund (ELBO) of the marginal likelihood for observed data, i.e., $P(X,A,Y)$, rather than optimizing the marginal likelihood directly. Specifically, we derive the negative ELBO, i.e., $\mathcal{L}_{\text{ELBO}}$, as follows:
@@
$+kl(q_{\phi_{1}}(Z_{A}|X,A)||p(Z_{A}))$
$+\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}(Z_{Y}|X,A)}\left[kl(q_{\phi_{2}}(\epsilon|X,A,Z_{Y})||p(\epsilon))\right]$
$+kl(q_{\phi_{3}}(Z_{Y}|X,A)||p(Z_{Y}))\vskip-17.22217pt$ (1)
@@
where $kl(\cdot||\cdot)$ denotes KL divergence. The derivation details are provided in Appendix A. $q_{\phi}$ indicates inference (encoder) network that approximates the posterior of latent variables, while $p_{\theta}$ indicates generative (decoder) network that models the likelihood of observed data given latent variables. Our objective is to find the optimal values of network parameters $\phi=$ $\phi_{1}$, $\phi_{2}$, $\phi_{3}$} and $\theta=$ $\theta_{1}$, $\theta_{2}$, $\theta_{3}$} that minimize the value of $\mathcal{L}_{\text{ELBO}}$. By doing so, the encoders and decoders are trained to directly capture the causal relationships among the variables that introduce noise. Consequently, it promotes the accurate inference of the latent clean node label $Z_{Y}$ and latent clean graph structure $Z_{A}$ to effectively perform the node classification and link prediction tasks even in the presence of DANG.
@@ 4.2 Model Instantiations
In this section, we present the details of the practical implementation and optimization of DA-GNN based on the learning objective, $\mathcal{L}_{\text{ELBO}}$. The overall architecture and detailed algorithm of DA-GNN are provided in Fig 3 and Algorithm 1 in Appendix, respectively. The key challenge of the instantiation is how to accurately infer the latent variables $Z_{A}$, $Z_{Y}$, and $\epsilon$ in the presence of noisy $X$, AA, and $Y$. To alleviate the challenge, we design the robust inference encoders (Fig 3(a)) and generative decoders (Fig 3(b)) with the corresponding regularizers (Fig 3(a)) and reconstruction losses (Fig 3(b)). Consequently, the encoders would be able to accurately infer the latent variables by capturing the causal relationships among the variables that introduce noise.
@@ 4.2.1 Modeling Inference Encoder
Modeling $q_{\phi_{1}}(Z_{A}|X,A)$. The objective of modeling $q_{\phi_{1}}(Z_{A}|X,A)$ is to accurately infer the latent clean graph structure $Z_{A}$ that enhances the message passing of a GNN model. We obtain the latent graph $\mathbf{\hat{A}}=\{\hat{p}_{ij}\}_{N\times N}$, where $\hat{p}_{ij}=\rho(s(\mathbf{Z}_{i},\mathbf{Z}_{j}))$ and $\mathbf{Z}=\text{GCN}_{\phi_{1}}(\mathbf{X},\mathbf{A})$, and regularize $\mathbf{\hat{A}}$ based on the prior knowledge that pairs of nodes with high $\gamma$-hop subgraph similarity are more likely to form assortative edges zhao2023graphglow ; consm ; rsgnn , thereby encouraging $\mathbf{\hat{A}}$ to predominantly include such edges. This regularization is equivalent to minimizing $kl(q_{\phi_{1}}(Z_{A}|X,A)||p(Z_{A}))$ in Eqn. 1. However, computing $\hat{p}_{ij}$ in every epoch is impractical for large graphs, i.e., $O(N^{2})$. To this end, we pre-define a proxy graph based on the subgraph similarity, and compute $\hat{p}_{ij}$ as edge weights on the proxy graph. Please refer to the Appendix B for detailed information on implementation details.
Modeling $q_{\phi_{3}}(Z_{Y}|X,A)$. The objective of modeling $q_{\phi_{3}}(Z_{Y}|X,A)$ is to accurately infer the latent clean node label $Z_{Y}$. To this end, we instantiate the encoder $\phi_{3}$ as a GCN classifier. Specifically, we infer $Z_{Y}$ through $\hat{\mathbf{Y}}=\text{GCN}_{\phi_{3}}(\mathbf{X},\mathbf{\hat{A}})\in\mathbb{R}^{N\times C}$. We introduce the node classification loss ${\color[rgb]{.75,0,.25}\mathcal{L}_{\text{cls-enc}}}=\sum_{i\in\mathcal{V}^{L}}\text{CE}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{i})$, where CE is the cross entropy loss. To further enhance the quality of inference of $Z_{Y}$, we regularize $Z_{Y}$ to satisfy class homophily mcpherson2001birds by minimizing the KL divergence between the probability predictions $\mathbf{\hat{Y}}$ of each node and its first order neighbors in $\mathbf{\hat{A}}$. The implemented loss function is given by:
${\color[rgb]{.75,0,.25}\mathcal{L}_{\text{hom}}}=\sum_{i\in\mathcal{V}}\frac{\sum_{j\in\mathcal{N}_{i}}\hat{p}_{ij}\cdot kl(\hat{\mathbf{Y}}_{j}||\hat{\mathbf{Y}}_{i})}{\sum_{j\in\mathcal{N}_{i}}\hat{p}_{ij}},$ (2)
where $\mathcal{N}_{i}$ denotes the set of first-order neighbors of node $v_{i}$ within $\hat{\mathbf{A}}$. It is worth noting that this regularization is equivalent to minimizing $kl(q_{\phi_{3}}(Z_{Y}|X,A)||p(Z_{Y}))$ in Eqn. 1.
@@
The objective of modeling $q_{\phi_{22}}(\epsilon_{A}|X,A)$ is to accurately infer the structure noise incurring variable $\epsilon_{A}$ that determines whether each edge is clean or noisy. To this end, we regard $\epsilon_{A}$ as a set of scores indicating the likelihood of each observed edge being clean or noisy. To estimate the likelihood, we utilize small loss approach han2018co . Precisely, we compute the set of link prediction losses as $\{(1-\hat{p}^{el}_{ij})^{2}|(i,j)\in\mathcal{E}\}$, where $\hat{p}^{el}_{ij}$ represents the $\hat{p}_{ij}$ value at the final epoch during early-learning phase. Therefore, an edge with high $\hat{p}^{el}_{ij}$ value can be considered as a clean edge, and we instantiate $\epsilon_{A}$ as $\{\hat{p}_{ij}^{el}|(i,j)\in\mathcal{E}\}$.
To alleviate the uncertainty of a single training point’s loss value, we adopt an exponential moving average (EMA) technique: $\hat{p}_{ij}^{el}\leftarrow\xi\hat{p}_{ij}^{el}+(1-\xi)\hat{p}_{ij}^{c}$, where $\hat{p}_{ij}^{c}$ indicates the value of $\hat{p}_{ij}$ at the current training point, and $\xi$ indicates the decaying coefficient fixed to 0.9. This approach is equivalent to minimizing $kl(q_{\phi_{22}}(\epsilon_{A}|X,A)|p(\epsilon_{A}))$, where $p(\epsilon_{A})$ is assumed to follow the same distribution as $q_{\phi_{22}}(\epsilon_{A}|X,A)$ but with lower variance.
For the encoder $\phi_{21}$, we use an MLP that takes $X$ and $Z_{Y}$ as inputs and infers $\epsilon_{X}$. Additionally, we regularize $p(\epsilon_{X})$ to follow the standard multivariate normal distribution, which means that a closed form solution of $kl(q_{\phi_{21}}(\epsilon_{X}|X,Z_{Y})||p(\epsilon_{X}))$ can be obtained as ${\color[rgb]{.75,0,.25}\mathcal{L}_{\text{p}}}=-\frac{1}{2}\sum_{j=1}^{d_{2}}(1+\log\sigma_{j}^{2}-\mu_{j}^{2}-\sigma_{j}^{2})$ kingma2013auto , where $d_{2}$ is the dimension of a $\epsilon_{X}$. Note that these two regularization techniques are equivalent to minimizing $\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[kl(q_{\phi_{2}}(\epsilon|X,A,Z_{Y})||p(\epsilon))\right]$ in Eqn. 1.